OpenStudy (icalibear):

sqrt 3a+ 1 = a-3 Solve for a

OpenStudy (icalibear):

\[\sqrt{3a+1}= a - 3\]

OpenStudy (egbeach):

how do you get rid of a square root?

OpenStudy (icalibear):

square root both sides?

OpenStudy (icalibear):

i mean square both sides

OpenStudy (egbeach):

correct. what do you get when you do that?

OpenStudy (icalibear):

\[3a + 1 = a^2 + 9\]

OpenStudy (egbeach):

you should actually get \[3a+1=\left( a-3 \right)^{2}\] you want to square it all together

OpenStudy (icalibear):

oh okay, and then FOIL that?

OpenStudy (egbeach):

yep what do you get after you FOIL?

OpenStudy (icalibear):

\[(a - 3)(a-3)\] \[a^2 - 3a - 3a + 9\] \[a^2 - 6 + 9\]

OpenStudy (icalibear):

6a not 6

OpenStudy (egbeach):

okay good. so now the whole equation is \[3a+1= a ^{2}-6a+9\] Now set the eqation to 0

OpenStudy (icalibear):

\[a^2 - 6a + 9 - 3a - 1 = 0\]

OpenStudy (egbeach):

simplfy

OpenStudy (icalibear):

\[a^2 - 9a + 10 = 0\]

OpenStudy (egbeach):

9-1+8 not 10

OpenStudy (icalibear):

oh i thought it wa 9 - (-1)

OpenStudy (egbeach):

= 8*

OpenStudy (icalibear):

\[a^2 - 9a + 8 = 0\]

OpenStudy (egbeach):

yep. can you factor it out by yourself or do you need help?

OpenStudy (icalibear):

do i use an 8 and a 1

OpenStudy (egbeach):

yep!

OpenStudy (icalibear):

\[(a-8)(a+1)\] \[(a - 8) = 0\] \[a = 8\] \[(a+1) = 0\] \[a =-1\]

OpenStudy (egbeach):

it should be (a-1)=0 so both answers will be positive

OpenStudy (icalibear):

oh okay. thanks

OpenStudy (egbeach):

welcome